# Quantum Quasinormal Mode Theory for Dissipative Nano-Optics and Magnetodielectric Cavity Quantum Electrodynamics

### v2

@@ Abstract
The unprecedented pace of evolution in nanoscale architectures for cavity quantum electrodynamics (cQED) has posed crucial challenges for theory, where the quantum dynamics arising from the non-perturbative dressing of matter by cavity electric and magnetic fields, as well as the fundamentally non-hermitian character of the system are to be treated without significant approximation. The lossy electromagnetic resonances of photonic, plasmonic or magnonic nanostructures are described as quasinormal modes (QNMs), whose properties and interactions with quantum emitters and spin qubits are central to the understanding of dissipative nano-optics and magnetodielectric cQED. Despite recent advancements toward a fully quantum framework for QNMs, a general and universally accepted approach to QNM quantization for arbitrary linear media remains elusive. In this work, we introduce a unified theoretical framework, based on macroscopic QED and complex coordinate transformations, that achieves QNM quantization for a wide class of spatially inhomogeneous, dissipative (with possible gain components) and dispersive, linear, magnetodielectric resonators. The complex coordinate transformations equivalently convert the radiative losses into non-radiative material dissipation, and via a suitable transformation that reflects all the losses of the resonator, we define creation and annihilation operators that allow the construction of modal Fock states for the joint excitations of field-dressed matter. By directly addressing the intricacies of modal loss in a fully quantum theory of magnetodielectric cQED, our approach enables the exploration of modern, quantum nano-optical experiments utilizing dielectric, plasmonic, magnetic or hybrid cQED architectures, and paves the way towards a rigorous assessment of room-temperature quantum nanophotonic technologies without recourse to ad hoc quantization schemes.
@@ I Introduction
Cavity-mediated control and enhancement of light-matter interaction has become a pervasive paradigm in modern quantum science and technology. Already decades in maturity, the use of micro- and nanoscale, optical, dielectric resonators to both harness and manipulate the photonic properties of semiconductor quantum dots has enabled the exploration of fundamental cavity quantum electrodynamics (cQED) in the solid state Vahala (2003); Khitrova et al. (2006); Kira and Koch (2011), as well as a host of device applications ranging from high-performance, non-classical light sources and spin-photon interfaces to ultrafast, all-optical switching for integrated photonic quantum computation and networks Hepp et al. (2019); Arakawa and Holmes (2020); Heindel et al. (2023). In stark contrast to their dielectric counterparts, plasmonic nanocavities offer the unique ability to confine light to extremely sub-wavelength scales and massively enhance local electromagnetic fields via their resonant surface plasmon modes, thereby constituting unique architectures for the exploration of light-matter interaction free from diffraction bounds Baumberg et al. (2019). Crucially, plasmonic resonators have ushered in an unprecedented era of room-temperature cQED at the nanoscale, facilitating access to the regime of strong coupling even at the profound limit of single molecular and quantum dot emitters under ambient conditions Chikkaraddy et al. (2016); Santhosh et al. (2016); Groß et al. (2018); Park et al. (2019), and opening up a compelling vista of potential applications in quantum nanotechnologies Xiong et al. (2021); Clarke and Hess (2024), extending from single-qubit coherent control Zhang et al. (2017) and ultrafast single-photon emission Hoang, Akselrod, and Mikkelsen (2016); Bogdanov et al. (2020); You et al. (2020) to near-field multipartite entanglement generation Bello et al. (2020); Bello, Kongsuwan, and Hess (2022) and quantum sensing Kongsuwan et al. (2019). In contrast to the electric dipole coupling between light and matter, cQED phenomena emerging from magnetic interactions have also been the subject of increasing attention. Leveraging collective microwave excitations of the spins in ferromagnetic and ferrimagnetic structures, the demonstration of such effects as photon-magnon and cavity-mediated magnon-qubit strong coupling Zhang et al. (2014); Hou and Liu (2019); Li et al. (2019); Neuman, Wang, and Narang (2020), have established cavity (nano-)magnonics as an attractive paradigm Rameshti et al. (2022), with interesting prospective applications in the engineering of synthetic gauge fields Gardin et al. (2024), cavity-mediated spin-photon interfaces at the nanoscale Chen et al. (2024), and the realization of quantum networks of entangled spins.
@@
As duly highlighted in the literature, the ability to describe the optical response of photonic, plasmonic or magnonic resonators and their interaction with quantum emitters (QEs) and spin qubits in terms of a small number of modes is of significant practical utility. A cavity mode formalism renders more transparent the physical attributes of light-matter interaction, allows one to isolate and quantify both radiative and non-radiative contributions to the observed dynamics, and aids device optimization towards particular applications. For ideal, closed cavities in the absence of material absorption (i.e., hermitian systems), normal modes emerge with real eigenfrequencies (resonances), infinite lifetimes and which are orthonormal with respect to some scalar product. In practice, the dielectric microcavities of traditional semiconductor quantum optics (which feature very high quality factors) most closely approximate this scenario Vahala (2003); Khitrova et al. (2006), but nevertheless, the notion remains ambiguous in general and cannot be applied without further considerations Kristensen, Vlack, and Hughes (2012); Kristensen and Hughes (2014). The issue of decay is especially important for nanoplasmonic resonators, where Ohmic dissipation inherent to their metallic constituents and open-cavity radiation losses are essentially universal, but has also garnered interest in the context of magnonic cavity systems where dissipative coupling effects have been observed Wang and Hu (2020); Yu et al. (2024) Of course, these losses typically have a deleterious impact on the performance of schemes for quantum information processing Xiong et al. (2021), but could also be harnessed in order to engineer the functionality of quantum metamaterials and to promote novel veins of research in non-hermitian topological photonics or in the quantum thermodynamic facets of nanophotonic technologies. In all such cases, it is evidently desirable to capture in full the inherently dissipative character of the electromagnetic resonances.
@@
To date, one of the most powerful and increasingly widespread approaches to this end has been quasinormal mode (QNM) theory Lai et al. (1990); Leung, Liu, and Young (1994); Leung et al. (1994); Leung and Pang (1996); Lee, Leung, and Pang (1999); Kristensen and Hughes (2014); Yan, Faggiani, and Lalanne (2018); Lalanne et al. (2018); Sauvan et al. (2022). The QNMs are time-harmonic solutions to Maxwell’s equations, satisfying suitable outgoing-wave boundary conditions (e.g., the Silver-Müller radiation condition) and bearing complex eigenfrequencies, loss thus being accounted for as an intrinsic modal property. The imposition of a radiation condition ensures that light propagates away from the cavity (as expected for a leaky resonator), but in conjunction with the complex nature of their eigenfrequencies, gives rise to a mathematically and conceptually difficult property of QNMs, namely their spatial, exponential divergence at large distances from the resonator. More generally, the non-conservative nature of lossy cavity systems has raised a variety of non-trivial questions regarding a proper, mathematically rigorous formulation of normalization and orthogonality, alongside the matter of whether or not the QNMs truly form a complete set, and thus allow mode expansions for arbitrary electromagnetic field distributions. In spite of the challenges and ambiguities, classical QNM methodology has undergone significant advances in the last decade, including improved understanding of the formal aspects of the theory Sauvan et al. (2022), techniques for computing the QNMs of arbitrary resonators Bai et al. (2013); Sauvan et al. (2013); Ge and Hughes (2014); Ge et al. (2014); Yan, Faggiani, and Lalanne (2018); Lalanne et al. (2018) and the release of general-purpose software Wu et al. (2023), as well as successful applications to the calculation of generalized effective mode volumes Kristensen, Vlack, and Hughes (2012); Sauvan et al. (2013); Kristensen and Hughes (2014), electromagnetic near- and far-field characteristics Kongsuwan et al. (2020); Ren et al. (2020); Bedingfield et al. (2023), Green’s functions Lee, Leung, and Pang (1999); Ge et al. (2014), enhanced spontaneous emission rates Sauvan et al. (2013); Ge et al. (2014); Kamandar Dezfouli, Gordon, and Hughes (2017); Ren et al. (2024) and even elastic Purcell factors in optomechanical systems El-Sayed and Hughes (2020).
@@
Inevitably, a quantum description of lossy electromagnetic modes is essential towards the fundamental study and understanding of quantum nanophotonic systems, including the statistics of photon, plasmon or magnon emission, polaritonic effects and the role of quantum fluctuations, as well as to facilitate rigorous quantification of key figures of merit for emerging quantum nano-optical technologies. To date, most theoretical investigations of nanoscale cQED and spectroscopy (e.g., surface-enhanced and electron energy loss spectroscopies) have relied on an ad hoc quantization protocol, in which the bosonic cavity mode is directly modeled as a quantum harmonic oscillator without formal, field-theoretic justification, while decay effects are incorporated via the relevant dissipators in a quantum master equation treatment, themselves derived from a phenomenological coupling to a bath of continuum modes Neuman, Aizpurua, and Esteban (2020); Giulio, Kociak, and de Abajo (2019); Zhou et al. (2022). Recently however, significant progress was made with the introduction of a second quantization scheme for QNMs, enabling the construction of multiplasmon or multiphoton Fock states for arbitrary, three-dimensional (3D), dissipative and non-magnetic resonators Franke et al. (2019). This approach has been elaborated and applied to a range of problems, such as concerning the limits of (phenomenological) dissipative Jaynes-Cummings models in plasmonic cQED, the performance of near-field-driven single-photon sources Hughes et al. (2019), the description of nonlinear cQED effects in hybrid metal-dielectric cavity systems Franke et al. (2020a) and coupled lossy and amplifying resonators Franke, Ren, and Hughes (2022). An important feature of these works is the regularization of the QNMs based on a Dyson equation formalism and a dual expansion scheme for the dyadic Green’s function Ge et al. (2014), which utilizes the unregularized modes for positions inside the resonator and regularized modes that are functions of real frequency for positions outside. Whilst the Dyson equation approach appears to yield an accurate representation for the Green’s function outside the resonator volume and allows accurate calculation of the Purcell factor in this region for the studied systems Ge et al. (2014), numerical computation of the regularized fields appears computationally intensive, and may even become intractable for more complex photonic cavity systems. Moreover, the formal justification for such a mode regularization technique remains elusive Sauvan et al. (2022), and in particular, is based on an assumed completeness of the QNMs inside the resonator. Whilst this may apply to simple geometries, (i.e., compact resonators in a homogeneous background medium), it is not guaranteed for more complex geometries, such as resonators on a substrate or waveguide-coupled resonator systems Sauvan et al. (2022). Alternative quantization schemes for lossy and non-magnetic resonators have also emerged in the last few years, relying on a fitting of the spectral density of the electromagnetic environment Medina et al. (2021) based on emitter-centered modes Feist, Fernández-Domínguez, and García-Vidal (2020), or a transformation of the continuum of photonic eigenmodes into a discrete set of pseudomodes Yuen and Demetriadou (2024), where the latter offers the notable advantage of circumventing traditional reservoir approximations and can thereby capture non-Markovian effects in the dynamics of nanocavity-QE systems. Nevertheless, it remains desirable to formulate a unified and numerically feasible methodology for describing the lossy resonances of arbitrary, linear, magnetodielectric systems, directly addressing such crucial issues as mode divergence, the identification of canonical field variables in the presence of both dielectric and magnetic media, and the calculation of key figures of merit for quantum nano-optical device applications on rigorous foundations. It is worth mentioning here that the quantization of QNMs and their coupling to QEs is similar to the reaction coordinate mapping of a (strongly-coupled) spin-boson model with a structured bath, described by a spectral density peaked around a certain frequency Binder et al. (2018); Anto-Sztrikacs and Segal (2021). In this method, the bosonic continuum is decomposed into a collective bosonic degree of freedom and a residual continuum, for which the interaction can be treated within the Markov approximation.
@@
In this work, we present a unified theoretical framework that achieves QNM quantization for a wide class of 3D, spatially inhomogeneous, dissipative (with possible gain components) and dispersive, linear, magnetodielectric resonators. Our approach is based on a rigorously defined mode regularization and truncation of the spatial domain via exterior complex coordinate transformations, tantamount to perfectly matched layers (PMLs) Hugonin and Lalanne (2005); Sauvan et al. (2022); Demésy et al. (2023). Harnessing established auxiliary-field eigenvalue techniques (or suitable generalizations thereof) with PMLs, a complete set of regularized modes can be ascertained, enabling the deployment of modal expansions for the dyadic Green’s function and electromagnetic fields both inside and outside the resonator volume. The quantization of these modes is then treated in the formalism of macroscopic QED (mQED), facilitating the construction of discrete Fock states for the joint excitations of field-dressed matter in a manner that fully accounts for both radiative and non-radiative modal losses, and which provides the foundations for a quantum dynamic theory via quantum Langevin and master equations. Our present methodology attains unprecedented scope among quantum QNM techniques, offering a consolidated approach to the analysis of contemporary quantum nano-optical experiments utilizing dielectric, plasmonic, magnonic or hybrid cQED architectures, and paves the way towards a rigorous and numerically feasible assessment of the performance metrics for emerging, room-temperature-viable, quantum nanophotonic technologies without recourse to ad hoc quantization schemes.
@@
The paper is organized as follows. § II briefly reviews the classical electromagnetic QNM theory and PML regularization, as well as the auxiliary-field eigenvalue technique that yields a complete set of orthonormalized modes, alongside useful formulae for time-domain analysis, in the commonly treated case of non-dispersive magnetic permeabilities. In § III, we discuss the theory of mQED in arbitrary, linear, magnetodielectric media; here, we quantize the modes by incorporating PML regularization in the formalism and defining bosonic modal operators in order to construct the corresponding Fock space. Using these operators, we derive an effective Lindblad master equation in § IV to describe the dynamics of the quantized modes and their coupling to QEs. In § V, we demonstrate the theory by investigating the spontaneous decay of a QE in a simple, 1D, half-open cavity, comparing our predictions with the semi-analytical results of an exact mQED treatment. Finally, we close the paper in § VI with some concluding remarks and perspectives.
@@
II Classical QNMs: Regularization, normalization, and completeness based on stretched coordinates
II.1 Modal formulation of the scattering problem

### v3

@@ Abstract
The unprecedented pace of evolution in nanoscale architectures for cavity quantum electrodynamics (cQED) has posed crucial challenges for theory, where the quantum dynamics arising from the non-perturbative dressing of matter by cavity electric and magnetic fields, as well as the fundamentally non-hermitian character of the system are to be treated without significant approximation. The lossy electromagnetic resonances of photonic, plasmonic or magnonic nanostructures are described as quasinormal modes (QNMs), whose properties and interactions with quantum emitters and spin qubits are central to the understanding of dissipative nano-optics and magnetodielectric cQED. Despite recent advancements toward a fully quantum framework for QNMs, a general and universally accepted approach to QNM quantization for arbitrary linear media remains elusive. In this work, we introduce a unified theoretical framework, based on macroscopic QED and complex coordinate transformations, that achieves QNM quantization for a wide class of spatially inhomogeneous, dissipative and dispersive, linear, magnetodielectric resonators. The complex coordinate transformations equivalently convert the radiative losses into non-radiative material dissipation, and via a suitable transformation that reflects all the losses of the resonator, we define creation and annihilation operators that allow the construction of modal Fock states for the joint excitations of field-dressed matter. By directly addressing the intricacies of modal loss in a fully quantum theory of magnetodielectric cQED, our approach enables the exploration of modern, quantum nano-optical experiments utilizing dielectric, plasmonic, magnetic or hybrid cQED architectures, and paves the way towards a rigorous assessment of room-temperature, quantum nanophotonic technologies without recourse to ad hoc quantization schemes.
@@ I Introduction
Cavity-mediated control and enhancement of light-matter interaction has become a pervasive paradigm in modern quantum science and technology. Already decades in maturity, the use of micro- and nanoscale, optical, dielectric resonators to both harness and manipulate the photonic properties of semiconductor quantum dots has enabled the exploration of fundamental cavity quantum electrodynamics (cQED) in the solid state vahala2003optical; khitrova2006vacuum; kirakochbook, as well as a host of device applications ranging from high-performance, non-classical light sources and spin-photon interfaces to ultrafast, all-optical switching for integrated photonic quantum computation and networks hepp2019qd; arakawaholmes2020progress; heindel2023qd. In stark contrast to their dielectric counterparts, plasmonic nanocavities offer the unique ability to confine light to extremely sub-wavelength scales and massively enhance local electromagnetic fields via their resonant surface plasmon modes, thereby constituting unique architectures for the exploration of light-matter interaction free from diffraction bounds baumberg2019extreme. Crucially, plasmonic resonators have ushered in an unprecedented era of room-temperature cQED at the nanoscale, facilitating access to the regime of strong coupling even at the profound limit of single molecular and quantum dot emitters under ambient conditions chikkaraddy2016single; santhosh2016vacuum; gross2018near; park2019tip, and opening up a compelling vista of potential applications in quantum nanotechnologies xiong2021room; clarke2024near, extending from single-qubit coherent control zhang2017sub and ultrafast single-photon emission hoang2016ultrafast; bogdanov2020ultrafast; you2020reconfigurable to near-field multipartite entanglement generation bello2020controlled; bello2022near and quantum sensing kongsuwan2019sensing. In contrast to the electric dipole coupling between light and matter, cQED phenomena emerging from magnetic interactions have also been the subject of increasing attention. Leveraging collective microwave excitations of the spins in ferromagnetic and ferrimagnetic structures, the demonstration of such effects as photon-magnon and cavity-mediated magnon-qubit strong coupling zhang2014strongly; hou2019strongmagnon; li2019magnon; neuman2020nanomagnonic, have established cavity (nano-)magnonics as an attractive paradigm rameshti2022cavity, with interesting prospective applications in the engineering of synthetic gauge fields gardin2024engineering, cavity-mediated spin-photon interfaces at the nanoscale chen2024scalable, and the realization of quantum networks of entangled spins.
@@
As duly highlighted in the literature, the ability to describe the optical response of photonic, plasmonic or magnonic resonators and their interaction with quantum emitters (QEs) and spin qubits in terms of a small number of modes is of significant practical utility. A cavity mode formalism renders more transparent the physical attributes of light-matter interaction, allows one to isolate and quantify both radiative and non-radiative contributions to the observed dynamics, and aids device optimization towards particular applications. For ideal, closed cavities in the absence of material absorption (i.e., hermitian systems), normal modes emerge with real eigenfrequencies (resonances), infinite lifetimes and which are orthonormal with respect to some scalar product. In practice, the dielectric microcavities of traditional semiconductor quantum optics (which feature very high quality factors) most closely approximate this scenario vahala2003optical; khitrova2006vacuum, but nevertheless, the notion remains ambiguous in general and cannot be applied without further considerations kristensen2012mode; kristensen2014modes. The issue of decay is especially important for nanoplasmonic resonators, where Ohmic dissipation inherent to their metallic constituents and open-cavity radiation losses are essentially universal, but has also garnered interest in the context of magnonic cavity systems where dissipative coupling effects have been observed wang2020dissipativemag; yu2024mag Of course, these losses typically have a deleterious impact on the performance of schemes for quantum information processing xiong2021room, but could also be harnessed in order to engineer the functionality of quantum metamaterials and to promote novel veins of research in non-hermitian topological photonics or in the quantum thermodynamic facets of nanophotonic technologies. In all such cases, it is evidently desirable to capture in full the inherently dissipative character of the electromagnetic resonances.
@@
To date, one of the most powerful and increasingly widespread approaches to this end has been quasinormal mode (QNM) theory lai1990time; leung1994completeness; leung1994perturbation; leung1996dielectric; lee1999dyadic; kristensen2014modes; yan2018rigorous; lalanne2018light; sauvan2022norm. The QNMs are time-harmonic solutions of Maxwell’s equations, satisfying suitable outgoing-wave boundary conditions (e.g., the Silver-Müller radiation condition) and bearing complex eigenfrequencies, loss thus being accounted for as an intrinsic modal property. The imposition of a radiation condition ensures that light propagates away from the cavity (as expected for a leaky resonator), but in conjunction with the complex nature of their eigenfrequencies, gives rise to a mathematically and conceptually difficult property of QNMs, namely their exponential divergence at large spatial distances from the resonator. More generally, the non-conservative nature of lossy cavity systems has raised a variety of non-trivial questions regarding a proper, mathematically rigorous formulation of normalization and orthogonality, alongside the matter of whether or not the QNMs truly form a complete set, and thus allow mode expansions for arbitrary electromagnetic field distributions. In spite of the challenges and ambiguities, classical QNM methodology has undergone significant advances in the last decade, including improved understanding of the formal aspects of the theory sauvan2022norm, techniques for computing the QNMs of arbitrary resonators bai2013efficient; sauvan2013theory; ge2014time; ge2014quasinormal; yan2018rigorous; lalanne2018light and the release of general-purpose software wu2023man, as well as successful applications to the calculation of generalized effective mode volumes kristensen2012mode; sauvan2013theory; kristensen2014modes, electromagnetic near- and far-field characteristics kongsuwan2020qnm; ren2020nfft; bedingfield2023multi, Green’s functions lee1999dyadic; ge2014quasinormal, enhanced spontaneous emission rates sauvan2013theory; ge2014quasinormal; dezfouli2017modal; ren2024classical, the semi-classical treatment of resonator-QE systems in the weak coupling regime yang2015analytical, and even elastic Purcell factors in optomechanical systems elsayed2020elastic.
@@
Inevitably, a quantum description of lossy electromagnetic modes is essential towards the fundamental study and understanding of quantum nanophotonic systems, including the statistics of photon, plasmon or magnon emission, polaritonic effects and the role of quantum fluctuations, as well as to facilitate rigorous quantification of key figures of merit for emerging quantum nano-optical technologies. To date, most theoretical investigations of nanoscale cQED and spectroscopy (e.g., surface-enhanced and electron energy loss spectroscopies) have relied on an ad hoc quantization protocol, in which the bosonic cavity mode is directly modeled as a quantum harmonic oscillator without formal, field-theoretic justification, while decay effects are incorporated via the relevant dissipators in a quantum master equation, themselves derived from a phenomenological coupling to a bath of continuum modes neuman2020sers; digiulio2019probing; zhou2022shb; hummer2013weak. Such a treatment is expected to work well if the resonator response is dominated by a single mode or if the spectral density has a simple, multi-Lorentzian profile in the multimode case, as in Ref. crookes2025collective for example, but breaks down if the spectral density exhibits a more complex structure. Recently however, significant progress was made with the introduction of a second quantization scheme for QNMs, enabling the construction of multiplasmon or multiphoton Fock states for arbitrary, three-dimensional (3D), dissipative and non-magnetic resonators franke2019quantization. This approach has been elaborated and applied to a range of problems, such as concerning the limits of (phenomenological) dissipative Jaynes-Cummings models in plasmonic cQED, the performance of near-field-driven single-photon sources hughes2019spe, the description of nonlinear cQED effects in hybrid metal-dielectric cavity systems franke2020nonlinear and coupled lossy and amplifying resonators franke2022coupled. An important feature of these works is the regularization of the QNMs based on a Dyson equation formalism and a dual expansion scheme for the dyadic Green’s function ge2014quasinormal, which utilizes the unregularized modes for positions inside the resonator and regularized modes that are functions of real frequency for positions outside. Whilst the Dyson equation approach appears to yield an accurate representation for the Green’s function outside the resonator volume and allows accurate calculation of the Purcell factor in this region for the studied systems ge2014quasinormal, numerical computation of the regularized fields appears computationally intensive, and may even become intractable for more complex photonic cavity systems. Moreover, the formal justification for such a mode regularization technique remains elusive sauvan2022norm, and in particular, is based on an assumed completeness of the QNMs inside the resonator. Whilst this may apply to simple geometries (i.e., compact resonators in a homogeneous background medium), it is not guaranteed for more complex geometries, such as resonators on a substrate or waveguide-coupled resonator systems sauvan2022norm. Furthermore, for non-trivial backgrounds, knowledge of the corresponding background Green’s function would be required for regularization, the determination of which could itself be non-trivial.
Alternative quantization schemes for lossy and non-magnetic resonators have also emerged in the last few years, relying on a fitting of the spectral density of the electromagnetic environment medina2021few; sanchez2022few based on emitter-centered modes buhmann2008casimir; feist2020macroscopic; hummer2013weak; rousseaux2016adiabatic, or a transformation of the continuum of photonic eigenmodes into a discrete set of pseudomodes yuen2024exact. The former offers a practical method to describe the interaction of QEs with the electromagnetic environment in terms of a few interacting lossy modes, and also requires knowledge of the photonic Green’s function. These modes, however, do not have a direct physical interpretation and their number is not fixed (scaling with the number of QEs featuring in the problem), while it remains unclear if the electric field can be correctly recovered from these discrete modes alone, as would be desirable for a modal theory. Nevertheless, fewer of these modes may be needed relative to the QNM case, which may become especially relevant in the vicinity of accumulation points of QNMs (e.g., due to high-wavevector plasmonic resonances). In contrast, the pseudomode theory offers the notable advantage of circumventing traditional reservoir approximations and can thereby capture non-Markovian effects in the dynamics of nanocavity-QE systems. Notably though, the theory is formulated on the basis of conventional QED, by assuming that the electric field can be expanded in terms of a continuum of normal modes, and consequently cannot be used to describe lossy or dispersive systems. Furthermore, it remains unclear how efficiently the method can be applied to systems that do not admit analytical tractability (in particular, for morphologically complex resonators), and if it can be used for arbitrary initial conditions beyond those pertinent to the single-excitation regime.
Notwithstanding the aforementioned efforts, it remains desirable to formulate a unified and numerically feasible methodology for describing the lossy resonances of arbitrary, linear, magnetodielectric systems, directly addressing such crucial issues as mode divergence, the identification of canonical field variables in the presence of both dielectric and magnetic media, and the calculation of key figures of merit for quantum nano-optical device applications on rigorous foundations. It is worth mentioning here that the quantization of QNMs and their coupling to QEs is similar to the reaction coordinate mapping of a (strongly-coupled) spin-boson model with a structured bath, described by a spectral density peaked around a certain frequency binder2018thermodynamics; anto2021capturing. In this method, the bosonic continuum is decomposed into a collective bosonic degree of freedom and a residual continuum, for which the interaction can be treated within the Markov approximation.
@@
In this work, we present a unified theoretical framework that achieves QNM quantization for a wide class of 3D, spatially inhomogeneous, dissipative (with possible gain components) and dispersive, linear, magnetodielectric resonators. Our approach is based on a rigorously defined mode regularization and truncation of the spatial domain via exterior complex coordinate transformations, tantamount to perfectly matched layers (PMLs) hugonin2005pml; sauvan2022norm; demesy2023dispersive. Harnessing established auxiliary-field eigenvalue techniques (or suitable generalizations thereof) with PMLs, a complete set of regularized modes can be ascertained, enabling the deployment of modal expansions for the dyadic Green’s function and electromagnetic fields both inside and outside the resonator volume. The quantization of these modes is then treated in the formalism of macroscopic QED (mQED), facilitating the construction of discrete Fock states for the joint excitations of field-dressed matter in a manner that fully accounts for both radiative and non-radiative modal losses, and which provides the foundations for a quantum dynamic theory via quantum Langevin and master equations. Our present methodology attains unprecedented scope among quantum QNM techniques, offering a consolidated approach to the analysis of contemporary quantum nano-optical experiments utilizing dielectric, plasmonic, magnonic or hybrid cQED architectures, and paves the way towards a rigorous and numerically feasible assessment of the performance metrics for emerging, room-temperature-viable, quantum nanophotonic technologies without recourse to ad hoc quantization schemes.
@@
The paper is organized as follows. § II briefly reviews the classical electromagnetic QNM theory and PML regularization, as well as the auxiliary-field eigenvalue technique that yields a complete set of orthonormalized modes, alongside useful formulae for time-domain analysis, in the commonly treated case of non-dispersive magnetic permeabilities. In § III, we discuss the theory of mQED in arbitrary, linear, magnetodielectric media; here, we quantize the modes by incorporating PML regularization in the formalism and defining bosonic modal operators in order to construct the corresponding Fock space. Using these operators, we derive an effective Lindblad master equation in § IV to describe the dynamics of the quantized modes and their coupling to QEs. In § V, we demonstrate the theory by investigating the spontaneous decay of a QE in a simple, 1D, half-open cavity (§ V.1) and in a 3D, spherical cavity (§ V.2), comparing our predictions with the semi-analytical results of an exact mQED treatment. Finally, we close the paper in § VI with some concluding remarks and perspectives.
@@
II Classical QNMs: Regularization, Normalization and Completeness Based on Stretched Coordinates
In this section, we review the classical electromagnetic QNM theory based on PML regularization, as introduced in Refs. sauvan2013theory; vial2014quasimodal; yan2018rigorous and reviewed in Ref. lalanne2018light. Our aim is to provide a reasonably self-contained presentation that captures the key concepts and important formulae that are also required in the quantized treatment, whilst additionally establishing the relevant notation.